1. Suppose you have a credit card with a 24.31% APR on purchases. You pay the minimum payment by the statement due date, leaving an outstanding balance of $3,200. How much would the balance be 1 day later? Assume that the card compounds daily.

a. $3,200.00 b. $3,264.56 c. $3202.13 d. $3,211.18

2. Suppose you have a credit card with a 24.31% APR on purchases. You pay the minimum payment by the statement due date, leaving an outstanding balance of $3,200. How much would the balance be 30 days later?

a. $3,200.00 b. $3202.13 c. $3,264.56 d. $3,211.18
Answer #1

1. 24.31%/365= .000666

3200*(1+.000666)= 3202.13.

option c is correct.

2. 24.31%/365*30=.019981

3200*(1+.019981)=3264

option c is correct